OpenStudy (anonymous):

the points are (-1,-4) and (2,1) Which equation is in point-slope form and depicts the equation of this line? (y+4)=−1/3(x+1) (y−1)=−1/3(x−2)(y−1)=−1/3(x−2) (y−4)=−5/3(x−1) (y+4)=5/3(x+1

OpenStudy (muzzack):

first find slope m=y2-y1/x2-x1 m=1-(-4)/2-(-1) m=5/3 then replace y-y1=m(x-x1) y+4=5/3x+1 so ur answer is D

OpenStudy (anonymous):

Why thank you.